# Spatially Resolved Kinematics of Post-starburst Galaxies

### v1

@@ Abstract
Star-forming galaxies can reach quiescence via rapid transition through merger-triggered starbursts that consequently affect both their kinematics and star formation rates. In this work, we analyze the spatially-resolved kinematics of 92 post-starburst galaxies (PSBs) with data from the MaNGA survey and place them in context with early-type galaxies (ETGs) to study the impact of merger history on galaxy kinematics. We measure the specific angular momentum to characterize them as slow or fast rotators. We find that the MaNGA PSB sample has $5\%$ slow rotators, which is less than the $14\%$ slow rotators in the ATLAS3D ETG sample and $22\%$ slow rotators in the MaNGA ETGs. The PSBs have lower specific angular momentum than the star-forming galaxies (SFs) and a greater fraction of slow rotators. This implies that for the PSBs to evolve into the ETGs, they must still lose some angular momentum. While ETGs with higher stellar mass tend to be slow rotators, PSBs do not follow this trend. We find significant correlations between specific angular momentum and mass-weighted age for the SF and ETG samples, but do not see any significant trends within the short PSB phase. These results indicate that significant evolution in angular momentum must continue to take place as the galaxy ages after the PSB phase. For PSBs to evolve directly into ETGs, they must undergo dry mergers to shed excess angular momentum without causing further epochs of star formation.
post-starburst galaxies — galaxy evolution — observations
software: Astropy (Astropy Collaboration et al., 2013, 2018), Matplotlib (Hunter, 2007), NumPy (Harris et al., 2020), SDSS-Marvin (Cherinka et al., 2019), Pandas (pandas development team, 2024), Seaborn (Waskom, 2021), SciPy (Virtanen et al., 2020)
@@ 1 Introduction
In order to trace the evolution of a galaxy, and any mergers undergone by it, we study its kinematics. Resolved integral field unit spectroscopy gives information about the rotational and dispersion components of the stellar velocity, which lets us quantify the specific angular momentum as done by Emsellem et al. (2007), and study the relative significance of stellar rotational velocity ($v$) and stellar velocity dispersion ($\sigma$). The specific angular momentum ($\lambda_{R}$) of a galaxy provides information about whether a galaxy is a slow ($\lambda_{R}\lesssim 0.1$) or a fast ($\lambda_{R}\gtrsim 0.1$) rotator (Emsellem et al., 2007). A comparison of this quantity for PSBs with other galaxies, especially ETGs, can help us connect the merger history with evolution in stellar populations.
@@
The angular momentum content of galaxies is sensitive to their recent merger history because the geometry of the merger and merger mass ratio will affect how much angular momentum is lost or gained. Classifying galaxies kinematically, based on $\lambda_{R}$, can offer valuable insight into their merger history. A loss in angular momentum can happen when two similar mass galaxies merge, in a major merger, along their centers such that they decrease the angular momentum of the resulting galaxy (Naab & Burkert, 2003). A gain in angular momentum could be a product of a merger between a massive galaxy colliding with a significantly lower mass galaxy (a minor merger) with a high value of angular momentum, which leads to the resulting galaxy gaining the excess momentum (Naab & Burkert, 2003). If the merger does not result in a huge influx of cold gas (a dry merger), it can lead to a loss in angular momentum without triggering further star formation.
@@
Observations of galaxies in the immediate aftermath of a strong starburst can be used to test the origins of angular momentum content in ETGs. Major mergers are thought to play a key role in the creation of ETGs (Hopkins et al., 2008), and thus, we would expect the ETGs to be slow-rotating and dispersion-dominated. However, Emsellem et al. (2007, 2011) found that ETGs consist of both slow and fast rotators. Emsellem et al. (2007) suggest that dissipationless (or “dry”) mergers, which cause angular momentum to be expelled outwards, play a key role in the evolution of slow rotators. They propose that multiple channels including dry mergers, gas accretion, and wet mergers likely play a role in the evolution of angular momentum in different classes of ETGs. A key test of how and when slow rotator ETGs lose their angular momentum is the fraction of slow rotators in galaxies in the immediate aftermath of a merger. PSB galaxies appear to be in the middle of evolving from the starburst phase to ETG phase, likely driven by a recent major merger between two gas-rich galaxies (Zabludoff et al., 1996; Yang et al., 2008; Wild et al., 2009; Pawlik et al., 2016; Sazonova et al., 2021), providing a snapshot into the ETG population that may result from this rapid channel of evolution.
@@
Norton et al. (2001), using long-slit spectroscopic data, found 18 out of 20 PSBs from Zabludoff et al. (1996) to be dispersion-dominated, i.e. $v/\sigma<1$, and only 6 out of 20 PSBs exhibited rotation with $v>40$ km/s. Pracy et al. (2009), Swinbank et al. (2012) and Pracy et al. (2013) used integral field spectroscopy to analyze the kinematics of PSBs and found that the PSBs are primarily dominated by fast rotators. Pracy et al. (2012) and Pracy et al. (2013) used the ellipticity-dependent classification presented in Emsellem et al. (2011) to find 4 slow rotators in a sample of 26 E+A galaxies at $z\sim 0.02-0.04$. They found that the slow rotator fraction for the PSBs ($\sim 15-17\%$) is similar to the slow rotator fraction in the ETGs from the ATLAS3D sample ($14\%$) (Emsellem et al., 2011). Similarly, Swinbank et al. (2012) found ${}9\%$ slow rotators in a sample of 11 E+A galaxies at $z=0.06-0.12$. In contrast to findings by Norton et al. (2001) that are more consistent with PSBs experiencing major mergers, results from Swinbank et al. (2012) and Pracy et al. (2013) suggest that their sample of PSBs probably experienced minor mergers. However, a lack of slow rotators in PSBs does not automatically discard the possibility of a major merger as even a 2:1 mass ratio merger can lead to a fast rotator (Emsellem et al., 2011) after triggering a starburst. Moreover, simulated slow rotators usually lose their angular momentum over a period of quite a few Gyrs via dry mergers (Naab et al., 2014). While it is clear that most PSBs are not already slow rotators, their dispersion-dominated kinematics and fraction of slow rotators are similar to ETGs within small number uncertainties. Larger samples of PSBs with spatially-resolved kinematics are needed to test how the angular momentum content of quiescent galaxies evolves over time, and the role of major and minor mergers in the formation of slow rotator ETGs.
@@
In order to further test and compare the results obtained by earlier works, we use integral field spectroscopic data for a large sample of 92 post-starburst galaxies from the Mapping Nearby Galaxies at APO (Apache Point Observatory) (MaNGA Bundy et al., 2015; Yan et al., 2016; Abdurro’uf et al., 2022) survey, which provides high-resolution spectroscopy for the inner kinematics of 10,000 nearby galaxies. We perform similar $\lambda_{R}$ calculations to Emsellem et al. (2007) for our PSB sample to test whether the PSBs are a result of major mergers, are evolving into ETGs or have undergone various minor/dry mergers. If all of the PSBs have been subjected to equal mass major mergers, we would expect them to have more slow rotators as compared to the ETGs, with perhaps another secular channel producing fast-rotator ETGs. If wet mergers are the main mechanism to produce ETGs, with slow rotators produced in the immediate aftermath of the merger-triggered starburst, we would expect PSBs and ETGs to have a comparable fraction of slow rotators. If minor dry mergers are required to produce slow rotator ETGs, we would expect to see fewer slow rotators among the PSB sample than in the ETG sample.
@@
We present the data and methods in §2. In §3.1, we outline the calculation and corrections performed for the specific angular momentum using the angular momentum expression given by Emsellem et al. (2007). We present our results from our analysis of angular momentum as a function of ellipticity, mass-weighted, light-weighted ages, and asymmetry in §3.2 and §4. Finally, we summarize our results and the implications from this work in §5.
@@
Figure 1: The figure shows the Marvin visualization of the MaNGA data for four of the identified slow rotators in the post-starburst galaxy sample. The left-most panel shows the SDSS gri postage stamp image of each galaxy, with the pink hexagon representing the MaNGA field of view, and the title gives the plate-ifu that can be used to identify the galaxy in Marvin and MaNGA. The middle and right-most panels show the corresponding stellar velocity and stellar velocity dispersion, each having a fixed color bar across the galaxies. The solid blue ellipse is created using the photometric properties from the NSA catalog.
@@
Figure 2: The figure shows the Marvin visualization of the MaNGA data for four of the identified fast rotators in the post-starburst galaxy sample. The left-most panel shows the SDSS gri postage stamp image of each galaxy, with the pink hexagon representing the MaNGA field of view, and the title gives the plate-ifu that can be used to identify the galaxy in MaNGA. The middle and right-most panels show the corresponding stellar velocity and stellar velocity dispersion, each having a fixed color bar across the galaxies. The solid blue ellipse is created using the photometric properties from the NSA catalog.
@@ 2.1 Post-starburst Galaxy Selection
There are many methods in the literature for selecting galaxies that have experienced a recent starburst (see French, 2021, for a recent review). This work selects galaxies with a variety of post-starburst ages and AGN properties from a combination of three different samples of PSBs identified in the Sloan Digital Sky Survey (SDSS, York et al. 2000; Strauss et al. 2002), as used in French et al. (2023).
@@ 2.2 MaNGA
The MaNGA Survey (Bundy et al., 2015; Yan et al., 2016; Abdurro’uf et al., 2022), is one of the three surveys of SDSS-IV that provides data collected for composition and kinematic structure of $\sim$10,000 nearby galaxies via the Sloan 2.5 m telescope at APO. It utilizes optical circular fibers to group 29 fiber bundles of various sizes to create hexagonally-shaped integral field units (IFUs) that are used to spectroscopically observe various sections of a galaxy for a wavelength range of $360-1040$ nm and a spectral resolution of 2000 (Drory et al., 2015; Law et al., 2015; Wake et al., 2017). The fibres have a core diameter of 120 microns which subtends to $2\arcsec$ on the sky. The Sloan telescope has 17 IFUs of various sizes that are used to observe different galaxies simultaneously within a field of view of $3$ deg diametrically ($7$ deg2).
We compared the right ascension and declination of the galaxies from the DR17 version of MaNGA to the $\sim$5000 galaxies from the combined post-starburst sample described in §2.1 to select 92 PSBs [^1] One galaxy is dropped from the sample used by French et al. (2023) because it is not included in the MPA-JHU galSpec analysis, which we use to obtain the stellar mass.. We used NASA-Sloan Atlas (NSA) catalog (Blanton et al., 2011) version v1_0_1 to obtain the photometric properties like Sérsic index (SERSIC_N), ellipticity ($\epsilon=1-$SERSIC_BA), position angle (SERSIC_PHI) and effective radius (SERSIC_TH50) in the r-band, all of which were calculated using a Sérsic fit, for the MaNGA galaxies. We explore the implication of choosing this photometry catalog in Appendix §A.
We obtain the light-weighted (Age_LW_Re_fit) and mass-weighted (Age_MW_Re_fit) ages, and star formation rates (log_SFR_ssp and log_SFR_Ha) for the MaNGA galaxies from SDSS Release 17 pyPipe3D pipeline (v3_1_1), which is a result of the analysis of stellar populations and ionized gas as presented in Sánchez et al. (2016a, b, 2018). Age_LW_Re_fit and Age_MW_Re_fit are the logarithms of the luminosity-weighted and mass-weighted ages of the stellar population in years within the effective radius of the galaxy, respectively. log_SFR_ssp is the logarithm of the integrated star formation rate obtained from the stellar mass that formed in the last 32 Myr. log_SFR_Ha is the logarithm of the integrated star formation rate determined from the integrated H-alpha flux (Sánchez et al., 2016b). We obtain the stellar mass from the “galSpec” galaxy properties from MPA-JHU (Kauffmann et al., 2003; Brinchmann et al., 2004; Tremonti et al., 2004). In order to distinguish the MaNGA SFs from the ETGs, we calculate the specific star-formation rate (sSFR) (SFR/M), where SFR is the integrated star formation rate obtained from the stellar mass log_SFR_ssp. A sSFR $\approx 10^{-11}$ yr-1 approximately divides star-forming and quiescent galaxies (Brinchmann et al., 2004). This sSFR cut is similar to the Halpha EW cut used by French et al. (2018) in selecting PSBs.
Graham et al. (2018) calculated and compared the angular momentum for 2300 galaxies of various types from MaNGA DR14 data, and identified the slow and fast rotators in their sample. We remove the PSBs we identified from their sample, as our angular momentum calculations for the PSBs are based on the MaNGA survey’s DR17 version. The data for the remaining MaNGA star-forming and early-type galaxies is utilized for comparisons and analysis in multiple upcoming plots.
@@
We use data products from the MaNGA data analysis pipeline (Westfall et al., 2019; Belfiore et al., 2019) to measure the angular momentum. We use marvin tools (Cherinka et al., 2019) to obtain the stellar velocity (stellar_vel), stellar velocity dispersion (stellar_sigma), and flux (spx_mflux) for each pixel given as defined by x-coordinate (spx_skycoo: on_sky_x) and y-coordinate (spx_skycoo: on_sky_y). These velocities have been offset to remove their cosmological redshift with respect to the solar barycentric rest frame. To perform this correction, stellar velocities returned by pPXF are used to calculate the observed redshift for individual spaxels, and the effect of the galaxy’s input cosmological redshift is removed to convert the velocities to the reference frame of the galaxy (Westfall et al., 2019). Similarly, the stellar velocity dispersion ($\sigma_{obs}$) is obtained by performing the pPXF fit for the template spectra to the MaNGA data. The $\sigma_{obs}$ needs to be corrected for the effective difference in the instrumental dispersion of the template spectra and MaNGA spectra (Westfall et al., 2019). This correction has been described in Eq 2.
@@
Example data products for the post-starburst galaxies are shown in Figures 1 and 2, which depict the observed galaxy images from SDSS (left-most panels), stellar velocity (middle panels), and stellar velocity dispersion (right-most panels). Figure 1 shows 4 slow rotators and Figure 2 shows 4 fast rotators, from the PSB sample, that were distinguished using the classification discussed in §3.2. These figures showcase the diversity in the post-starburst sample, as galaxies like 11752-3701 (slow rotator) and 9493-1902 (fast rotator), the top galaxies in each figure, have similar morphologies. However, they vary quite a bit in their kinematics, as depicted by the corresponding stellar velocity and dispersion plots. On the other hand, 12071-3702 (slow rotator) and 8080-3704 (fast rotator) are highly elliptical galaxies. The PSB sample also displays a range of merger signatures, like 9507-12704 (fast rotator), and tidal features, like 11753-3704 (fast rotator). These characteristics are quite evident in not just the SDSS images in the left panels, but also from the inner kinematics presented via stellar velocity and stellar velocity dispersion in the middle and right panels.
@@
Plateifu - Marvin Plate id (Abdurro’uf et al., 2022) to obtain spectroscopic data
MaNGA_Id - MaNGA Id (Abdurro’uf et al., 2022) for the galaxy
RA deg Right ascension (Aihara et al., 2011) of the center of the galaxy
Dec deg Declination (Aihara et al., 2011) of the center of the galaxy
Epsilon - Ellipticity (ϵ=italic-$\epsilon=$1 - SERSIC_BA (Blanton et al., 2011)) of the disk
FWHM_PSF arcsec Full-width half maximum (dispersion)222given by the point spread function which is assumed to be a gaussian (zfwhm)
Sersic_n - Sérsic index (SERSIC_N) (Blanton et al., 2011)
R_eff arcsec Half-light radius (SERSIC_TH50) (Blanton et al., 2011) of the galaxy
@@
Phot_angle deg Position angle (SERSIC_PHI) (Blanton et al., 2011) of the disk East of North
Lambda_Re - Corrected angular momentum at half-light radius
@@
SFR_ssp M/yr star formation rate (log_SFR_ssp) (Sánchez et al., 2016a, b, 2018) of the galaxy
SFR_ha M/yr star formation rate (log_SFR_Ha) (Sánchez et al., 2016a, b, 2018) of the galaxy in H-alpha
LW_age Gyr Light-weighted age (Age_LW_Re_fit) (Sánchez et al., 2016a, b, 2018) of the galaxy
MW_age Gyr Mass-weighted age (Age_MW_Re_fit) (Sánchez et al., 2016a, b, 2018) of the galaxy
Asymmetry - Morphological asymmetry (A) (Hernández-Toledo et al. (2010); Cheng et al. (2011); Vázquez-Mata et al. (2022))
of the galaxy
Table 1: PSB galaxy properties. Full table available online.

### v2

@@ Abstract
Star-forming galaxies can reach quiescence via rapid transition through merger-triggered starbursts that consequently affect both their kinematics and star formation rates. In this work, we analyze the spatially resolved kinematics of 89 post-starburst galaxies (PSBs) with data from the Mapping Nearby Galaxies at Apache Point Observatory (MaNGA) survey and place them in context with star-forming galaxies (SFGs) and early-type galaxies (ETGs) to study the impact of merger history on galaxy kinematics. We measure the specific angular momentum to characterize them as slow or fast rotators. We find that the MaNGA PSB sample has $\sim 6\%$ slow rotators, which is less than the $\sim 14\%$ slow rotators in the ATLAS3D ETG sample and $\sim 20\%$ slow rotators in the MaNGA ETGs, and more than the $\sim 3.5\%$ slow rotators in the MaNGA SFGs. This implies that for PSBs to evolve into ETGs, they must still lose some angular momentum. While ETGs with higher stellar mass tend to be slow rotators, PSBs do not follow this trend. We find significant correlations between specific angular momentum and mass-weighted age for the SFG and ETG samples, but do not see any significant trends within the short PSB phase. These results indicate that significant evolution in angular momentum must continue to take place as the galaxy ages after the PSB phase. For PSBs to evolve directly into ETGs, they must undergo dry mergers to shed excess angular momentum without causing further epochs of star formation.
E+A galaxies — Post-starburst galaxies — Galaxy kinematics — Galaxy evolution — Galaxy quenching
software: Astropy (Astropy Collaboration et al., 2013, 2018, 2022), Matplotlib (Hunter, 2007), NumPy (Harris et al., 2020), SDSS-Marvin (Cherinka et al., 2019), Pandas (pandas development team, 2024), Seaborn (Waskom, 2021), SciPy (Virtanen et al., 2020)
@@ 1 Introduction
In order to trace the evolution of a galaxy and any mergers undergone by it, we study its kinematics. Resolved integral field unit (IFU) spectroscopy gives information about the rotational and dispersion components of the stellar velocity, which lets us quantify the specific angular momentum, as done by Emsellem et al. (2007), and study the relative significance of stellar rotational velocity ($v$) and stellar velocity dispersion ($\sigma$). The specific angular momentum ($\lambda_{R}$) of a galaxy provides information about whether a galaxy is a slow ($\lambda_{R}\lesssim 0.1$) or a fast ($\lambda_{R}\gtrsim 0.1$) rotator (Emsellem et al., 2007). A comparison of this quantity for PSBs with other galaxies, especially ETGs, can help us connect the merger history with evolution in stellar populations.
@@
The angular momentum content of galaxies is sensitive to their recent merger history because the geometry of the merger and merger mass ratio will affect how much angular momentum is lost or gained. Classifying galaxies kinematically, based on $\lambda_{R}$, can offer valuable insight into their merger history. A loss in angular momentum can happen when two similar mass galaxies merge, in a major merger, along their centers such that they decrease the angular momentum of the resulting galaxy (Naab & Burkert, 2003). A gain in angular momentum could be a product of a merger between a massive galaxy colliding with a significantly lower-mass galaxy (a minor merger) with a high value of angular momentum, which leads to the resulting galaxy gaining the excess momentum (Naab & Burkert, 2003). If the merger does not result in a huge influx of cold gas (a dry merger), it can lead to a loss in angular momentum without triggering further star formation.
@@
Observations of galaxies in the immediate aftermath of a strong starburst can be used to test the origins of angular momentum content in ETGs. Major mergers are thought to play a key role in the creation of ETGs (Hopkins et al., 2008), and thus we would expect the ETGs to be slow-rotating and dispersion dominated. However, Emsellem et al. (2007, 2011) found that ETGs consist of both slow and fast rotators. Emsellem et al. (2007) suggest that dissipationless (or “dry”) mergers, which cause angular momentum to be expelled outwards, play a key role in the evolution of slow rotators. They propose that multiple channels including dry mergers, gas accretion, and wet mergers likely play a role in the evolution of angular momentum in different classes of ETGs. A key test of how and when slow-rotator ETGs lose their angular momentum is the fraction of slow rotators in galaxies in the immediate aftermath of a merger. PSB galaxies appear to be in the middle of evolving from the starburst phase to the early-type phase, likely driven by a recent major merger between two gas-rich galaxies (Zabludoff et al., 1996; Yang et al., 2008; Wild et al., 2009; Pawlik et al., 2016; Sazonova et al., 2021), providing a snapshot into the ETG population that may result from this rapid channel of evolution.
@@
Norton et al. (2001), using long-slit spectroscopic data, found 18 out of 20 PSBs from Zabludoff et al. (1996) to be dispersion dominated, i.e. $v/\sigma<1$, and only six out of 20 PSBs exhibited rotation with $v>40$ km s-1. Pracy et al. (2009), Swinbank et al. (2012) and Pracy et al. (2013) used integral field spectroscopy to analyze the kinematics of PSBs and found that the PSBs are primarily dominated by fast rotators. Pracy et al. (2012) and Pracy et al. (2013) used the ellipticity-dependent classification presented in Emsellem et al. (2011) to find four slow rotators in a sample of 26 E+A galaxies at $z\sim 0.02-0.04$. They found that the slow-rotator fraction for the PSBs ($\sim 15\%-17\%$) is similar to the slow-rotator fraction in the ETGs from the ATLAS3D sample ($14\%$; Emsellem et al. (2011)). Similarly, Swinbank et al. (2012) found $\sim 9\%$ slow rotators in a sample of 11 E+A galaxies at $z=0.06-0.12$. In contrast to findings by Norton et al. (2001) that are more consistent with PSBs experiencing major mergers, results from Swinbank et al. (2012) and Pracy et al. (2013) suggest that their sample of PSBs probably experienced minor mergers. However, a lack of slow rotators in PSBs does not automatically discard the possibility of a major merger as even a 2:1 mass ratio merger can lead to a fast rotator (Emsellem et al., 2011) after triggering a starburst. Moreover, simulated slow rotators usually lose their angular momentum over a period of quite a few billion years via dry mergers (Naab et al., 2014). While it is clear that most PSBs are not already slow rotators, their dispersion-dominated kinematics and fraction of slow rotators are similar to ETGs within small number uncertainties. Larger samples of PSBs with spatially resolved kinematics are needed to test how the angular momentum content of quiescent galaxies evolves over time, and the role of major and minor mergers in the formation of slow-rotator ETGs.
@@
In order to further test and compare the results obtained by earlier works, we use integral field spectroscopic data for a large sample of 89 PSBs from the Mapping Nearby Galaxies at Apache Point Observatory (MaNGA; Bundy et al., 2015; Yan et al., 2016; Abdurro’uf et al., 2022) survey, which provides high-resolution spectroscopy for the inner kinematics of 10,000 nearby galaxies. We perform similar $\lambda_{R}$ calculations to Emsellem et al. (2007) for our PSB sample to test whether the PSBs are a result of major mergers, are evolving into ETGs or have undergone various minor/dry mergers. If all of the PSBs have been subjected to equal-mass major mergers, we would expect them to have more slow rotators as compared to the ETGs, with perhaps another secular channel producing fast-rotator ETGs. If wet mergers are the main mechanism to produce ETGs, with slow rotators produced in the immediate aftermath of the merger-triggered starburst, we would expect PSBs and ETGs to have a comparable fraction of slow rotators. If minor dry mergers are required to produce slow-rotator ETGs, we would expect to see fewer slow rotators among the PSB sample than in the ETG sample.
@@
We present the data and methods in Section 2. In Section 3.1, we outline the calculation and corrections performed for the specific angular momentum using the angular momentum expression given by Emsellem et al. (2007). We present our results from our analysis of angular momentum as a function of ellipticity, mass-weighted (MW) age, light-weighted (LW) age, and asymmetry in Sections 3.2 and 4. Finally, we summarize our results and the implications from this work in Section 5.
@@
Figure 1: The Marvin visualization of the MaNGA data for four of the identified slow rotators in the post-starburst galaxy sample. The left-most panel shows the SDSS gri postage stamp image of each galaxy, with the pink hexagon representing the MaNGA field of view, and the title gives the plate-ifu that can be used to identify the galaxy in Marvin and MaNGA. The middle and rightmost panels show the corresponding stellar velocity and stellar velocity dispersion, each having a fixed color bar across the galaxies. The solid blue ellipse is created using the photometric properties from the NSA catalog.
@@
Figure 2: The Marvin visualization of the MaNGA data for four of the identified fast rotators in the post-starburst galaxy sample. The left-most panel shows the SDSS gri postage stamp image of each galaxy, with the pink hexagon representing the MaNGA field of view, and the title gives the plate-ifu that can be used to identify the galaxy in MaNGA. The middle and rightmost panels show the corresponding stellar velocity and stellar velocity dispersion, each having a fixed color bar across the galaxies. The solid blue ellipse is created using the photometric properties from the NSA catalog.
@@ 2.1 Post-starburst Galaxy Selection
There are many methods in the literature for selecting galaxies that have experienced a recent starburst (see French, 2021, for a recent review). This work selects galaxies with a variety of post-starburst ages and active galactic nuclei (AGN) properties from a combination of three different samples of PSBs identified in the Sloan Digital Sky Survey (SDSS; York et al. 2000; Strauss et al. 2002), as used in French et al. (2023).
@@ 2.2 MaNGA
The MaNGA Survey (Bundy et al., 2015; Yan et al., 2016; Abdurro’uf et al., 2022) is one of the three surveys of SDSS-IV that provides data collected for the composition and kinematic structure of $\sim$10,000 nearby galaxies via the Sloan 2.5 m telescope at the Apache Point Observatory. It utilizes optical circular fibers to group 29 fiber bundles of various sizes to create hexagonally shaped IFUs that are used to spectroscopically observe various sections of a galaxy for a wavelength range of $360-1040$ nm and a spectral resolution of 2000 (Drory et al., 2015; Law et al., 2015; Wake et al., 2017). The fibres have a core diameter of 120 $\mu$m which subtends to $2\arcsec$ on the sky. The Sloan telescope has 17 IFUs of various sizes that are used to observe different galaxies simultaneously within a field of view of $3^{\circ}$ diametrically ($7$ deg2).
We compared the R.A. and decl. of the galaxies from the Data Release 17 (DR17) version of MaNGA to the $\sim$5000 galaxies from the combined post-starburst sample described in Section 2.1 to select 89 PSBs [^1] One galaxy is dropped from the sample used by French et al. (2023) because it is not included in the MPA-JHU galSpec analysis, which we use to obtain the stellar mass. Three galaxies were dropped as they are undergoing mergers (based on their stellar velocity maps), which can significantly impact their angular momentum, thus introducing inaccuracies in kinematic analysis. This leaves us with 89 PSBs out of the 93 in French et al. (2023).. We used NASA-Sloan Atlas (NSA) catalog (Blanton et al., 2011) version v1_0_1 to obtain the photometric properties like Sérsic index (SERSIC_N), ellipticity ($\epsilon=1-$SERSIC_BA), position angle (SERSIC_PHI) and effective radius (SERSIC_TH50) in the $r$ band, all of which were calculated using a Sérsic fit, for the MaNGA galaxies. Graham et al. (2018) calculated and compared the angular momentum for 2300 galaxies of various types from MaNGA Data Release 14 (DR14) data, and identified the slow and fast rotators in their sample. We remove the 16 PSBs we identified from their sample, as our angular momentum calculations for the PSBs are based on the MaNGA survey’s DR17 version. The remaining galaxies (SFGs and ETGs) are cross-matched with the NSA catalog to reproduce their angular momentum calculation using MaNGA DR17 and NSA photometry. These MaNGA galaxies are utilized for comparisons and analysis in multiple upcoming plots. We prefer MaNGA release DR17 over DR14 as the sophisticated spectral fitting using enhanced stellar population models in DR17 results in more accurate measurements of emission lines and kinematic properties, and more accurate ages and metallicities. The DR17 data-analysis pipeline (DAP) has been updated to include better calibrated kinematic maps and tailored analysis options, which allow for finer analysis of kinematic properties (see Law et al., 2021, for more details). Most importantly, $\sim 10,000$ galaxies from DR17 include 89 PSBs as compared to the 16 PSBs identified in the $\sim 2800$ galaxies present in DR14, allowing us to perform a more robust comparison between the various samples. We explore the implication of choosing NSA photometry and DR17 instead of DR14 in Appendix A.
[…]
@@
We use data products from the MaNGA data-analysis pipeline (Westfall et al., 2019; Belfiore et al., 2019) to measure the angular momentum for the MaNGA SFGs, PSBs and ETGs. We use Marvin tools (Cherinka et al., 2019) to obtain the stellar velocity (stellar_vel), stellar velocity dispersion (stellar_sigma), and flux (spx_mflux) for each pixel given as defined by x-coordinate (spx_skycoo: on_sky_x) and y-coordinate (spx_skycoo: on_sky_y). These velocities have been offset to remove their cosmological redshift with respect to the solar barycentric rest frame. To perform this correction, stellar velocities returned by pPXF are used to calculate the observed redshift for individual spaxels, and the effect of the galaxy’s input cosmological redshift is removed to convert the velocities to the reference frame of the galaxy (Westfall et al., 2019). Similarly, the stellar velocity dispersion ($\sigma_{obs}$) is obtained by performing the pPXF fit for the template spectra to the MaNGA data. The $\sigma_{obs}$ needs to be corrected for the effective difference in the instrumental dispersion of the template spectra and MaNGA spectra (Westfall et al., 2019). This correction has been described in Equation (2).
@@
Example data products for the PSBs are shown in Figures 1 and 2, which depict the observed galaxy images from SDSS (leftmost panels), the stellar velocity (middle panels), and the stellar velocity dispersion (rightmost panels). Figure 1 shows four slow rotators and Figure 2 shows four fast rotators from the PSB sample that were distinguished using the classification discussed in Section 3.2. These figures showcase the diversity in the PSB sample, as galaxies like 11752-3701 (slow rotator) and 9493-1902 (fast rotator), the top galaxies in each figure, have similar morphologies. However, they vary quite a bit in their kinematics, as depicted by the corresponding stellar velocity and dispersion plots. On the other hand, 12071-3702 (slow rotator) and 8080-3704 (fast rotator) are highly elliptical galaxies, and galaxies like 7960-12704 (fast rotator) are disk-like and have spiral arms. The PSB sample also displays a range of merger signatures and tidal features, like 11753-3704 (fast rotator). These characteristics are quite evident in not just the SDSS images in the left panels, but also from the inner kinematics presented via the stellar velocity and stellar velocity dispersion in the middle and right panels.
Table 1: MaNGA galaxy properties [^3] This table is available in its entirety in machine-readable form in the online article.
@@
Plateifu Marvin Plate ID to obtain spectroscopic data (Abdurro’uf et al., 2022)
MaNGA_Id MaNGA ID for the galaxy (Abdurro’uf et al., 2022)
RA deg R.A. of the center of the galaxy (Aihara et al., 2011)
Dec deg Decl. of the center of the galaxy (Aihara et al., 2011)
Epsilon Ellipticity (ϵ=italic-$\epsilon=$1 - SERSIC_BA) of the disk (Blanton et al., 2011)
FWHM_PSF arcsec Full-width at half-maximum (dispersion) [^4] Given by the point-spread function, which is assumed to be a Gaussian (zfwhm)
Sérsic_n Sérsic index (SERSIC_N) (Blanton et al., 2011)
R_eff arcsec Half-light radius (SERSIC_TH50) of the galaxy (Blanton et al., 2011)
@@
Phot_angle deg Position angle (SERSIC_PHI) of the disk east of north (Blanton et al., 2011)
Lambda_Re Corrected angular momentum at half-light radius
@@
SFR_ssp M yr-1 Star formation rate (log_SFR_ssp) of the galaxy (Sánchez et al., 2016a, b, 2018)
[…]
LW_age Gyr Luminosity-weighted age (Age_LW_Re_fit) of the galaxy (Sánchez et al., 2016a, b, 2018)
MW_age Gyr Mass-weighted age (Age_MW_Re_fit) of the galaxy (Sánchez et al., 2016a, b, 2018)
Asymmetry Morphological asymmetry (A) of the galaxy
(Hernández-Toledo et al., 2010; Cheng et al., 2011; Vázquez-Mata et al., 2022)
Type Type of galaxy (SFG, PSB or ETG)